OpenStudy (anonymous):

If 3^x^2 / 3^4x = 1/81, then x =

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

i think x=1

OpenStudy (anonymous):

3^x^2? I don't get this

OpenStudy (anonymous):

no it is wrong

OpenStudy (anonymous):

(3)^ squared x

OpenStudy (anonymous):

\[\sqrt[3]{x} ?\]

OpenStudy (anonymous):

\[\sqrt[x]{3}\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

at the bottom equation rewrite the equation

OpenStudy (anonymous):

3^(x)^2

OpenStudy (anonymous):

\[\sqrt{3x}\]

OpenStudy (anonymous):

they don't write like you they write like I told you

OpenStudy (anonymous):

\[\sqrt{3^x}\]

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I don't get, you in Viet Nam now?

OpenStudy (anonymous):

no usa

OpenStudy (anonymous):

you don't understand what I write?

OpenStudy (anonymous):

no, look at the bottom equation rewrite the equation

OpenStudy (anonymous):

^ what you call ?

OpenStudy (anonymous):

is it square?

OpenStudy (anonymous):

square=()^2

OpenStudy (anonymous):

2^2

OpenStudy (anonymous):

\[3x^2\]

OpenStudy (anonymous):

so 3^x

OpenStudy (anonymous):

\[3^x\]

OpenStudy (anonymous):

and X^2

OpenStudy (anonymous):

x^2=\[x^2\]

OpenStudy (anonymous):

double ^

OpenStudy (anonymous):

triple^ ^ raise to power

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^3, x^4 , x\[x^3,x^4,X^5\]^5

OpenStudy (anonymous):

I know but it 3^x and x has ^2

OpenStudy (anonymous):

you understand now?

OpenStudy (anonymous):

\[3^{x^2}\]

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

exactly now

OpenStudy (anonymous):

\[\frac{3^{x^2}}{3^{4x}}=\frac{1}{81}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cancle? mean

OpenStudy (anonymous):

wait I type wrong let me start over again

OpenStudy (anonymous):

\[3^{x^2-4x}=3^{-4}\]\[x^2-4x+4=0\]\[(x-2)(x-2)=0\]or\[(x-2)^2=0\]x=2

OpenStudy (anonymous):

you right

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

are you Vietnamese?

OpenStudy (anonymous):

yes you can see your and my last name

OpenStudy (anonymous):

my husband the same your last name

OpenStudy (anonymous):

which state you live?

OpenStudy (anonymous):

California

OpenStudy (anonymous):

Medal help?

OpenStudy (anonymous):

n your?

OpenStudy (anonymous):

wow nice to meet u

OpenStudy (anonymous):

u too

OpenStudy (anonymous):

indiana

OpenStudy (anonymous):

u in Indian?

OpenStudy (anonymous):

what u study now?

OpenStudy (anonymous):

no I am vietnamese like u

OpenStudy (anonymous):

I live the north of USa

OpenStudy (anonymous):

indiana

OpenStudy (anonymous):

I,know , indiana what state?I never heard

OpenStudy (anonymous):

u in usa but u still use VN name

OpenStudy (anonymous):

I never hear indiana state?

OpenStudy (anonymous):

north of USA

OpenStudy (anonymous):

my name is Tham baker

OpenStudy (anonymous):

I,m mother , I go college to learn math try help my son math , your name don't sound in usa

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

Can u speak VN

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

good, my kid not at all

OpenStudy (anonymous):

how old are u?

OpenStudy (anonymous):

ok thank for ur help

OpenStudy (anonymous):

any way ,good see you , I will see you later, I have to go ,u in high school? I'm old over 40 ,sory i'm one the phone , don't forget give me madel,I will see you next time math help

OpenStudy (anonymous):

ok thanks